OpenStudy (anonymous):

6.03] The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today? 900 620 520 450

OpenStudy (anonymous):

a + b = 1070 3a - 2b = 110 solve for a and b where "a" is the # of one-step equations and "b" is the # of two-step equations.

OpenStudy (anonymous):

I suggest you use substitution or elimination.

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

idk what to doo

OpenStudy (anonymous):

Take your first equation, a + b = 1070 Solve for "a", which is just subtracting "b" from each side Put this value of "a" into the second equation.. That equation will have only one variable, "b". Solve for "b". Go back to your first equation with that value of "b" and get "a".

OpenStudy (anonymous):

i know that but how do i get the specific number

OpenStudy (anonymous):

That IS the way to get your specific number. There is only one value for "a" and one value for "b" that will satisfy the two equations. Just follow my steps.

OpenStudy (anonymous):

620?

OpenStudy (anonymous):

"a" is one-step and "b" is two-step. You gave the answer for the # of two-step equations. Your problem asks for the # of one-step equations. You're almost there with your answer. You're going in the right direction, though.

OpenStudy (anonymous):

its 520

OpenStudy (anonymous):

Check to see if you have the right answer by substituting those values for "a" and "b" back into the equations in my first post.